2. A certain radioactive substance has a half-life of 38 hours. Find how long it takes for 90% of

the radioactivity to be dissipated.


1
Expert's answer
2022-03-29T08:27:45-0400

"x = x_oe^{-kt}"

When t = 38 hr, x = 0.5xo

"0.5x_o = x_oe^{-38k}"

"e^{-k} = 0.5^{1\/38}"

Hence,

"x = x_o(0.5^{t\/38})"

When 90% are dissipated, x = 0.1x0

"0.1x_o = x_o(0.5^{t\/38})"

"0.1^{38} = 0.5^t"

"t = \dfrac{38 \ln 0.1}{\ln 0.5}"

"t = 126.23 ~ \text{hrs}"